THE SPECTROSCOPIC CONSEQUENCES OF C-H⋯π HYDROGEN BONDING: THE C_{6}H_{6}-C_{4}H_{2} COM-PLEX

\underline{TALITHA M. SELBY}, PAUL R. WINTER, CHRISTOPHER RAMOS, JAIME A. STEARNS, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-2084.

Resonant two-photon ionization spectroscopy and resonant ion-dip infrared spectroscopy are used to structurally characterize the C_{6}H_{6}-C_{4}H_{2} complex. The complex takes on a T-shaped geometry in which the diacetylene molecule sits on the six-fold axis of benzene, with one of its C-H groups hydrogen bonded to the π cloud of benzene. The direction and magnitude of the electronic frequency shift of the complex’s transitions relative to benzene monomer, the forbidden S_{0}-S_{1} origin transition, and the acetylenic C-H stretch vibrational fundamentals are all consistent with this structure. The complexation with benzene localizes the C-H stretch vibrations of the two acetylenic C-H groups. Both C-H stretch fundamentals are clearly observed, one unshifted and one shifted from its frequency in C_{4}H_{2} monomer. The C-H stretch fundamental of the hydrogen bonded C-H group is shifted down in frequency by about 45 cm^{−1} and gains significant intensity, as one would expect from a hydrogen bonded XH group. This complex is a good candidate for half-collision reaction dynamics following excitation of the C_{4}H_{2} molecule in the complex.